Please help me with combined andlike terms. The question is...-9(6+u)-2u=-10
Thank you, Chris

Answer by stanbon(75887)

-9(6+u)-2u=-10
---
-54 - 9u - 2u = -10
----
-11u = -10 + 54
-11u = 44
---
u = -4
==============
Cheers,
Stan H.
==============